# Domain-Specific Code Language Models: Unraveling the Potential for HPC Codes and Tasks

### v1

@@ Abstract
With easier access to powerful compute resources, there is a growing trend in AI for software development to develop larger language models (LLMs) to address a variety of programming tasks. Even LLMs applied to tasks from the high-performance computing (HPC) domain are huge in size and demand expensive compute resources for training. This is partly because these LLMs for HPC tasks are obtained by finetuning existing LLMs that support several natural and/or programming languages. We found this design choice confusing — why do we need large LMs trained on natural languages and programming languages unrelated to HPC for HPC-specific tasks?
In this line of work, we aim to question choices made by existing LLMs by developing smaller LMs for specific domains — we call them domain-specific LMs. Specifically, we start with HPC as a domain and build an HPC-specific LM, named MonoCoder, which is orders of magnitude smaller than existing LMs but delivers similar, if not better performance, on non-HPC and HPC tasks. Specifically, we pre-trained MonoCoder on an HPC-specific dataset (named HPCorpus) of C and C++ programs mined from GitHub. We evaluated the performance of MonoCoder against conventional multi-lingual LLMs. Results demonstrate that MonoCoder, although much smaller than existing LMs, achieves similar results on normalized-perplexity tests and much better ones in CodeBLEU competence for high-performance and parallel code generations. Furthermore, fine-tuning the base model for the specific task of parallel code generation (OpenMP parallel for pragmas) demonstrates outstanding results compared to GPT, especially when local misleading semantics are removed by our novel pre-processor Tokompiler, showcasing the ability of domain-specific models to assist in HPC-relevant tasks.
The sources of this work are available at our GitHub MonoCoder repository.
@@ I Introduction
Recent breakthroughs in the field of AI have led significant attention to language models (LMs) due to their remarkable capabilities in natural language processing [4] and understanding. Large language models (LLMs) [5], particularly exemplified by models such as GPT-3 [6] and its successors [7], have demonstrated the potential to grasp intricate linguistic structure and context, sparking exploration of their applicability beyond natural language tasks. Recently, GPT-3.5 and GPT-4 have been deployed in conversational environments in the form of ChatGPT [8].
@@
In parallel, the field of high-performance computing (HPC) has been tackling increasingly complex and data-intensive problems [9]. The field of HPC has experienced notable advancements in hardware, software, and algorithms, resulting in substantial improvements in computational performance and efficiency [10, 11]. Combining the two trends, integrating LMs into HPC workflows has emerged as a compelling avenue for innovation [12]. For instance, several recent efforts have applied LLMs for automatically inserting OpenMP pragmas or MPI functions in code [13, 14, 15, 16, 17, 18], overcoming limitations of static tools [19, 20, 21, 22, 23].
Although existing LLMs have shown remarkable results on programming tasks [24, 25], such as code generation, bug fixing, etc., to our surprise, we found that they performed poorly on some of HPC-related programming tasks such as code parallelization, vectorization, etc. [26, 15, 16, 12]. This led us to ponder on several research questions: (i) How well do existing LLMs perform on domain-specific tasks such as those for HPC domain, (ii) Will more domain-specific training data help in improving the performance?, (iii) When applying LMs for a specific domain, do we finetune existing LMs, or should we train them from scratch? This last question is specifically important in the context of training costs. For instance, HPC-Coder [16], a recently-introduced LM for HPC tasks, is obtained by finetuning PolyCoder [27] on an HPC dataset; PolyCoder itself is a code LM (not specific to HPC) that is trained on 249 GB of code corpus that is made up of programs written in 12 programming languages. HPC-GPT [28], LM4HPC [12], and AutoParLLM [29] rely on LLaMA as their base model — with up to 34 billion parameters — or even on GPT-4. Such setups looked counter-intuitive to us — Is it not enough to train an LM on HPC-specific languages only? In other words, why do we need an LM trained on Java or Python — with tens or hundreds of billion parameters — for HPC-specific tasks? More importantly, we believe that such domain-specific LMs would be computationally as well as financially efficient to train.
In this paper, we hypothesize that HPC-specific LMs (e.g., smaller LMs that are designed and trained specifically on HPC datasets) would perform better than existing LMs for HPC tasks. We perform two experiments to validate this hypothesis. The purpose of the first experiment is to build a smaller LM that performs equally well as existing LLMs on generic perplexity score. Towards that end, in the first experiment, we build a smaller LM, called MonoCoder, by reducing the number of layers of PolyCoder code LM by a factor of 4 and pretraining on merely C and C++ codes. We then empirically validate that MonoCoder, despite being a smaller model than PolyCoder, achieves a comparable perplexity score to PolyCoder on the HPCorpus dataset (HPC-specific dataset of C and C++ programs obtained from GitHub). The purpose of the second experiment is then to evaluate the performance of MonoCoder on HPC-specific tasks. Towards that end, in the second experiment, we begin with CodeBLEU competence for high-performance and parallel code generations given increasing context. Subsequently, we finetune MonoCoder and other existing LMs on the dataset of OpenMP parallelization task – pairs of for loops and corresponding OpenMP pragmas. We then evaluate the ability of these models to predict OpenMP private and reduction clauses accurately, including their corresponding variables. We found that existing code LMs capture “local” semantics of code structures (e.g., variable named i is an index variable of a for loop), which leads to their degraded performance on this test. We address this limitation in MonoCoder by developing a novel code pre-processing scheme called Tokompiler, which eliminates any local semantics that the LLM may capture. We found that MonoCoder outperforms existing code LMs in virtually all those tests and tasks, with and without applying Tokompiler. Moreover, and in contrast to MonoCoder, the other LMs’ performances considerably degraded when Tokompiler was applied, suggesting their limited understanding under semantic-less and uncommon settings.
Contributions. We make the following contributions in the paper:
  • By drawing insights from the limitations of existing LLMs, we propose a domain-specific model called MonoCoder for solving tasks related to high-performance computing.
  • We suggest a pre-processing method, called Tokompiler, for evaluating models’ code structure understanding by eliminating local semantics.
  • We compare the proposed approach against PolyCoder and GPT-3.5 — state-of-the-art LLMs — for tasks related to HPC programming.
  • We evaluate MonoCoder on (1) CodeBLEU competence for high-performance and parallel code generations given increasing context and the downstream task of (2) automatic shared memory parallelization via OpenMP, that highlights the strength of the domain-specific model in solving HPC-related tasks.
The remainder of the paper is organized as follows: In section II, we provide motivating examples that underscore the need to build domain-specific LMs. In section III, we unfold the key aspects of HPC programming and how they differ from general programming. Following this, section IV delves into the details of the HPC code dataset HPCorpus and preprocessing steps that should maximize code understanding from the LM. Then section V provides an in-depth exploration of MonoCoder, our language model pre-trained on HPCorpus. section VI introduces Tokompiler, our novel code pre-processor designed for HPC code processing in LMs, detailing its unique approach to generating anonymizing code snippets based on Abstract Syntax Trees (AST). Finally, in section VII, we perform evaluations of MonoCoder and state-of-the-art models on non-HPC and HPC tasks, while in section VIII, we present our efforts in fine-tuning MonoCoder for a specific HPC task, OpenMP pragma generation, demonstrating the practical implications of our work. We discuss related work in section IX. Finally, we conclude the paper in section X and outline directions for future research in section XI.
II Motivations for Building Domain-specific LMs
Recent work has proposed that domain-specific LMs trump general-purpose ones for many use cases. For example, [30] recently showed that a smaller LM trained solely on StackOverflow data outperforms larger models on comprehension tasks such as answer quality prediction and obsolete post detection. In non-software domains, others have shown that domain-specific pretraining can benefit comprehension tasks in biomedicine [31] and finance [32]. It stands to reason that domain-specific pre-training of code LMs may benefit HPC tasks. Here we discuss two specific examples that highlight the limitations of existing general-purpose code LMs.
(a) Correct answer by ChatGPT
(b) Incorrect answer by ChatGPT
Figure 1: ChatGPT example of OpenMP Parallelization Task: “Can this code be parallelized”.
Evaluating ability of GPT-3.5 on an HPC task. Given the popularity of ChatGPT and its ability to assist in programming tasks, we decided to experiment with the underlying GPT-3.5 model [^1] We could not use GPT-4 — an improved version of GPT-3.5 — as it was a premium- and paid-only version. via ChatGPT on the task of parallelizing simple for loops written in C language [15, 26]. We found that (0(a)), ChatGPT could correctly determine the parallelization potential of simple loops (with trivial access patterns) and also provide correct explanations of its determinations [^2] Because GPT-3.5 could parallelize simpler loops with correct explanations, it is reasonable to conclude that its training data included examples of parallel programs written in C language. But when we presented a trickier loop with a non-trivial access pattern (as shown in 0(b)), ChatGPT parallelized the loop incorrectly while also presenting an explanation that was contrary to the prior explanation.
Evaluating ability of StarCoder on a code completion task. Given that GPT-3.5 was designed to solve various tasks, in this experiment, we decided to use an LLM that we specifically developed and trained for programming tasks. Towards that end, we decided to use StarCoder [33], which has caught significant attention in the programming community recently. StarCoder models are 15.5B parameter models pretrained on 80+ programming languages from GitHub (1T training tokens).
Figure 2: An example showing code completion output of StarCoder. Black text shows the input prompt, blue is the generated completion.
In this experiment (Figure 2), we asked the StarCoder model via the HuggingFace website API [^3] https://huggingface.co/bigcode/starcoder to complete a simple C code by providing int main (int argc, const char *argv[]) as the starting prompt. We expected the model to produce an example of “Hello World” written in C or C++. However, to our surprise, it produced strange C code, followed by Python code! We can give a point to the model for producing C code (even though weird-looking), but it is difficult to comprehend why the model produced Python code. Further exploration revealed that the StarCoder model is finetuned on Python from the pretrained, multilingual StarCoderBase model. Naturally, StarCoder has a strong inclination towards Python, which explains why it produced Python code when given C code. Nevertheless, the creators claim that StarCoder “maintains strong performance on other programming languages.” [^4] https://huggingface.co/spaces/bigcode/bigcode-playground. Overall, we believe this example demonstrates that pretrained, multi-lingual models fine-tuned to other datasets can still get confused about the intended language of the input code.
These examples underscore the need for domain-specific training on particular tasks (as opposed to training on several tasks) with a limited number of input languages. This is precisely what we aimed to achieve in this exploration.
III Characteristics of General vs. HPC Programmers
The roles of general programmers and HPC programmers diverge significantly in terms of their tasks, optimization priorities and techniques, tool sets, and development environments [28, 12]. These distinctions arise from the unique demands of HPC programming, where computational speed and efficiency are paramount. In this section, we briefly talk about specific differences that are important in the context of building LLMs.
  1. 1.
    Nature of Tasks: General programmers typically engage in a diverse array of software development tasks, ranging from web and mobile application development to database management and scripting. In contrast, HPC programmers specialize in addressing highly complex scientific and computational problems. Their domain of expertise encompasses tasks such as weather forecasting [34], fluid dynamics simulations [35], structural analysis, and molecular modeling. HPC programmers are deeply involved in the development of applications that demand intensive mathematical modeling and parallel processing capabilities.
  2. 2.
    Performance Optimization: General programmers typically prioritize aspects like code readability, maintainability, and the expeditious development of software. While they acknowledge the importance of performance, it often takes a secondary role in their development efforts. Conversely, HPC programmers place performance at the forefront of their concerns. They meticulously optimize their code to achieve the utmost computational speed. This optimization entails parallelizing code, fine-tuning algorithms, and tailoring code for hardware architectures.
  3. 3.
    Tools and Libraries: General programmers work with a wide spectrum of programming languages and libraries, selecting them based on project requirements. Common programming languages in this context include Python, JavaScript, Java, and C#. In contrast, HPC programmers employ specialized tools and libraries, uniquely suited for scientific and parallel computing tasks. These encompass Message Passing Interface (MPI) [36] for distributed computing, OpenMP [37] for shared-memory parallelism, Intel’s HPC toolkit [38], and CUDA for GPU programming [39]. Furthermore, HPC programmers often work with domain-specific libraries tailored for scientific simulations, with C and C++ being the dominant programming languages.
  4. 4.
    Parallelism and Concurrency: General programming often involves the creation of single-threaded or multi-threaded applications, with parallelism not always being the central concern. Developers may employ threading or asynchronous programming to manage concurrency. HPC programming, conversely, is fundamentally rooted in parallelism. HPC programmers design their applications from the ground up to efficiently leverage parallel processing capabilities [38]. They commonly collaborate with clusters [40], supercomputers, and GPUs to harness maximum computational power.
  5. 5.
    Performance Profiling and Optimization: General programmers employ profiling tools to identify performance bottlenecks in their code. Their optimization efforts may include algorithmic improvements, code refactoring, or selecting appropriate libraries. HPC programmers, on the other hand, rely on advanced profiling and optimization tools such as Intel VTune [41] or NVIDIA Nsight [42] to scrutinize code performance at a granular level. They may engage in comprehensive code rewrites to minimize communication overhead, maximize vectorization, and optimize memory access patterns.
IV HPCorpus: HPC Code Corpus
@@
[…]
@@
V MonoCoder: An HPC-specific code LM
MonoCoder. To create a domain-specific model for HPC, we pre-trained a decoder-only transformer model on the language modeling objective (i.e., given past tokens as context, predict the next token in the code) using only C and C++ code from HPCorpus. We excluded the Fortran portion of the dataset given its relatively small size (Table I) and its overall dissimilarity to C/C++. We named this domain-specific model MonoCoder, in contrast to the multilingual Polycoder.
Model size. In the pursuit of deploying domain-specific small language models on average computer systems, careful consideration of model size becomes paramount. To ensure compatibility with such resource constraints, we aimed to design a model that strikes a balance between complexity and memory efficiency. The choice of a model comprising just under 1B parameters emerged as an optimal compromise. The decision is underpinned by the following calculations: Assuming each parameter is represented as a 32-bit floating-point number requiring 4 bytes, the formula RAM size (in bytes) = Number of Parameters $\times$× Size of one parameter (in bytes). Consequently, for 0.9B parameters, the resulting RAM size amounted to 3.6GB. This judiciously selected model size not only accommodates a 4GB RAM constraint – the maximum memory per core in HPC systems [45] – but also leaves additional headroom for other computational processes within the system to accommodate model inference. Note that this worst-case analysis did not consider inference-time optimizations such as quantization, pruning, etc., which would reduce memory consumption further.
Limitations. Our domain-specific model inherits the principles of a Left-to-Right Language Model, which is especially amenable to code generation tasks. By adhering to the left-to-right nature, our model aligns with the established models such as CodeGPT (124M) [46], CodeParrot (1.5B) [47], GPT-Neo [48] and PolyCoder (2.7B) [27], GPT-J (6B) [49], Codex (12B) [50], StarCoder (15.5B) [33], and GPT-NeoX (20B) [51], known for their proficiency in code completion.
However, we acknowledge the inherent challenge posed by the left-to-right approach, which limits the model’s ability to consider context beyond the immediate token sequence. As part of our ongoing research, we are exploring strategies to address this limitation and further refine MonoCoder’s capability to leverage broader contextual information, thereby advancing its effectiveness in the nuanced landscape of HPC code synthesis. This will be demonstrated especially in the design of our downstream tasks in section VIII.
Figure 3: MonoCoder training and validation loss curves.
Pre-training details. We pre-trained the MonoCoder model using the GPTNeoX framework [52] on 4 NVIDIA A40 48GB GPUs with fp16 precision. For pre-training, we used the Adam optimizer and followed a learning rate schedule with a linear warmup for the first 1% of steps and a cosine decay over the remaining steps. Gradients were clipped at 1.0. Each training sample had a maximum length of 2048 tokens and was trained in mini-batches of 16 samples (4 per GPU). The model was trained for 160K steps at a learning rate of 0.00008. Figure 3 shows the training and validation loss curves of MonoCoder.
VI Tokompiler: Preprocessing to Evaluate Code Structure Understanding
(a) GPT3 Tokenizer
(b) Codex Tokenizer
(c) GPT-3.5 & 4
(d) Tokompiler
Figure 4: GPT3, Codex, GPT-3.5&4 and Tokompiler tokenization of the Pi C code. As models are enhanced, tokenization is also improved, but it is still imperfect for domain-specific language primitives and numbers.
// Source code:
int main() {
  int r[2800 + 1];
// Tokompiler:
int func_252() {
 int arr_88[num_34 +
            num_842];
@@ VI Tokompiler: Preprocessing to Evaluate Code Structure Understanding
Figure 5: Tokompiler pipeline overview: Given a source code, the code turns into a semantic-less version using AST knowledge, and eventually, the lexicalized tokens are fed into MonoCoder.

### v3

# MonoCoder: Domain-Specific Code Language Model for HPC Codes and Tasks

@@ Abstract
With easier access to powerful compute resources, there is a growing trend in AI for software development to develop large language models (LLMs) to address a variety of programming tasks. Even LLMs applied to tasks from the high-performance computing (HPC) domain are huge in size and demand expensive compute resources for training. This is partly because LLMs for HPC tasks are obtained by finetuning existing LLMs that support several natural and/or programming languages. We found this design choice confusing — why do we need LLMs trained on natural languages and programming languages unrelated to HPC for HPC-specific tasks?
In this line of work, we aim to question choices made by existing LLMs by developing smaller language models (LMs) for specific domains — we call them domain-specific LMs. Specifically, we start with HPC as a domain and build an HPC-specific LM, named MonoCoder, which is orders of magnitude smaller than existing LMs but delivers better performance on non-HPC and HPC codes. Specifically, we pre-trained MonoCoder on an HPC-specific dataset (named HPCorpus) of C and C++ programs mined from GitHub. We evaluated the performance of MonoCoder against state-of-the-art multi-lingual LLMs. Results demonstrate that MonoCoder, although much smaller than existing LMs, outperforms other LLMs on normalized-perplexity tests (in relation to model size) while also delivering competing CodeBLEU scores for high-performance and parallel code generations. In other words, results suggest that MonoCoder understands HPC code better than state-of-the-art LLMs.
MonoCoder source code is available at our GitHub repository.
@@ I Introduction
Recent breakthroughs in the field of AI have led significant attention to language models (LMs) due to their advanced capabilities in natural language processing (NLP) [1]. Large language models (LLMs), particularly exemplified by models such as GPT-3 [2] and its successors [3], deployed in the conversational form of ChatGPT [4], have demonstrated the potential to grasp intricate linguistic structure and semantics, sparking exploration of their applicability beyond NLP.
@@
In parallel, the field of high-performance computing (HPC) has been tackling increasingly complex and data-intensive problems [5]. The field of HPC has experienced advancements in hardware, software, and algorithms, resulting in improvements in computational performance and efficiency [6, 7]. Combining the two trends, integrating LMs into HPC workflows has emerged as a compelling avenue for innovation [8]. For instance, several recent efforts have explored the application of LLMs in assisting HPC programmers in automatically inserting OpenMP pragmas or MPI functions in code [9, 10, 11, 12, 13, 14, 15, 16], overcoming the limitations of existing static tools [17, 18, 19, 20, 21].
@@
This finding led us to ponder several research questions: (i) How well do existing LLMs perform on domain-specific tasks such as those for the HPC domain, (ii) Will more domain-specific training data help in improving the performance?, (iii) When applying LMs for a specific domain, do we finetune existing LMs, or should we train them from scratch? This last question is specifically important in the context of the enormous training costs of LLMs, which is the second limitation of existing LLMs. As an example, HPC-Coder [13], a recently-introduced LM for HPC tasks, is obtained by finetuning PolyCoder [24] on an HPC dataset; PolyCoder itself is a code LM (not specific to HPC) that is trained on a corpus made up of 249 GB of programs written in 12 programming languages. As another example, HPC-GPT [25], LM4HPC [8], and AutoParLLM [26] rely on LLaMA as their base model — with up to 34 billion parameters — or even on GPT-4. Such setups looked counter-intuitive to us — Is it not enough to train an LM on HPC-specific languages only? In other words, why do we need an LM trained on Java or Python — with tens or hundreds of billion parameters — for HPC-specific tasks? More importantly, we believe that such domain-specific LMs would be computationally as well as financially efficient to train.
In this paper, we hypothesize that HPC-specific LMs (e.g., smaller LMs that are designed and trained specifically on HPC datasets) would perform better than existing LMs for HPC tasks. We perform two experiments to validate this hypothesis. The purpose of the first experiment is to build a smaller LM that performs similar, if not better, than existing LLMs in terms of generic language understanding tasks (such as code completion). Towards that end, in the first experiment, we build a smaller LM, called MonoCoder, by reducing the number of layers of PolyCoder by a factor of 4 and pretraining only on C and C++ codes. We then empirically validate that MonoCoder, despite being a smaller model than PolyCoder, achieves a comparable perplexity score to PolyCoder in generic code completion task. The purpose of the second experiment is then to evaluate the performance of MonoCoder on HPC-specific tasks. Towards that end, we obtain the CodeBLEU scores of parallel code generation models with increasing context.
We exploit one insight that we learned while building MonoCoder. Specifically, we found that existing code LMs capture “local” semantics of code structures that leads to their degraded performance on the code completion task. An example of “local” semantics could be a variable named i is an index variable of a for loop (in most of the programs for that matter). We address this limitation in MonoCoder by implementing a code pre-processing scheme that eliminates any local semantics that the LLM may capture. Our experimental evaluation demonstrates that MonoCoder outperforms existing code LMs in virtually all of the code completion settings and context lengths, with and without applying local semantics elimination (henceforth, LSE). Moreover, and in contrast to MonoCoder, the performance of other LMs degraded considerably when LSE was applied, suggesting their reliance on local semantics.
Contributions. This paper makes following contributions:
  • By drawing insights from the limitations of existing LLMs, we propose a domain-specific, small language model, called MonoCoder, for tasks related to high-performance computing.
  • We design a pre-processing method, called Local Semantics Elimination (LSE), that eliminates syntactical constructs that could lead to local semantics.
  • We compare MonoCoder against PolyCoder and GPT-3.5 for general-purpose programming tasks as well as tasks related to HPC programming.
  • Finally, we also measure the parallel code generation performance of MonoCoder and other LLMs over a dataset of 20k OpenMP codes by calculating the CodeBLEU score. Our results demonstrate that MonoCoder, although orders of magnitude smaller in size, performs similar to these LLMs in language comprehension, while outperforming them in HPC-specific tasks.
The remainder of the paper is organized as follows: In section II we delve into the details of the HPC code dataset HPCorpus. Then section III provides an in-depth exploration of MonoCoder, our language model pre-trained on HPCorpus. In section IV, we introduce LSE, our code preprocessing step designed for HPC code. In section V, we perform evaluations of MonoCoder and state-of-the-art models on non-HPC and HPC tasks. Finally, we conclude the paper and outline directions for the future research in section VII.
@@
1// Source code:
2int main() {
3 int r[2800 + 1];
4}
1// LSE:
2int func_252() {
3 int arr_88[num_34 + num_842];
4}
@@
Figure 1: Local Semantics Elimination (LSE) pipeline overview: Given a source code, the code turns into a semantic-less version using AST knowledge, and eventually, the lexicalized tokens are fed into MonoCoder.
II HPCorpus: HPC Code Corpus
In order to build an HPC-specific model, we first decided to gather a dataset of HPC specific programs. Towards that end, we compiled HPCorpus, a dataset of publicly-visible C and C++, and Fortran programs from GitHub [27]. In this work we decided to focus on C and C++ languages only, and hence we used a subset of HPCorpus as shown in Table I.
In [27], we discovered that many of those repos employed one (or more) parallel API. For instance, 45% employed shared memory parallelism with OpenMP (primarily for threading, but also for SIMD and GPU offloading), 27% employed distribution with MPI, 21% employed direct GPU programming (half with CUDA and half with OpenCL), and the rest mainly employed other threading APIs (such as TBB and Cilk).
[…]
III MonoCoder: An HPC-specific code LM
To create a domain-specific model for HPC, we pre-trained a decoder-only transformer model on the language modeling objective (i.e., given past tokens as context, predict the next token in the code) using only C and C++ code from HPCorpus. We named this domain-specific model as MonoCoder, in contrast to the multilingual Polycoder.
Model size. In the pursuit of deploying domain-specific small language models on average computer systems, careful consideration of model size becomes paramount. To ensure compatibility with such resource constraints, we aimed to design a model that strikes a balance between complexity and memory efficiency. The choice of a model comprising just under 1B parameters emerged as an optimal compromise. The decision is underpinned by the following calculations: Assuming each parameter is represented as a 32-bit floating-point number requiring 4 bytes, the formula RAM size (in bytes) = Number of Parameters $\times$× Size of one parameter (in bytes). Consequently, for 0.9B parameters, the resulting RAM size amounted to 3.6GB. This selected model size not only accommodates a 4GB RAM constraint – typically the maximum memory per core in HPC systems [28] – but also leaves additional headroom for other computational processes within the system to accommodate model inference. Note that this worst-case analysis did not consider inference-time optimizations such as quantization, or pruning, which would reduce memory consumption further.
Limitations. Our domain-specific model inherits the principles of a left-to-right language model, which is especially amenable to code generation tasks. By adhering to the left-to-right nature, our model aligns with the established models such as CodeGPT (124M) [29], GPT-Neo [30] and PolyCoder (2.7B) [24], GPT-J (6B) [31], Codex (12B) [32], StarCoder (15.5B) [33], and GPT-NeoX (20B) [34], which are known for their proficiency in code completion.
However, we acknowledge the inherent challenge posed by the left-to-right approach, which limits the model’s ability to consider context beyond the immediate token sequence. As part of our ongoing research, we are exploring strategies to address this limitation and further refine MonoCoder’s capability to leverage broader contextual information, thereby advancing its effectiveness in the nuanced landscape of HPC code synthesis.
Pre-training details. We pre-trained the MonoCoder model using the GPTNeoX framework [35] on 4 NVIDIA A40 48GB GPUs with fp16 precision. For pre-training, we used the Adam optimizer and followed a learning rate schedule with a linear warmup for the first 1% of steps and a cosine decay over the remaining steps. Gradients were clipped at 1.0. Each training sample had a maximum length of 2048 tokens and was trained in mini-batches of 16 samples (4 per GPU). The model was trained for 160K steps at a learning rate of 0.00008. The training and validation losses initially start high at around 3.0, rapidly decrease over the first 25,000 steps, stabilize thereafter, converge closely without overfitting, and finally settle below 0.5 by approximately 150,000 steps, indicating minimal loss.
MonoCoderPolyCoderGPT-NeoGPT-JCodexStarCoderGPT-NeoX11115555101010101515151520202020252525250.90.90.90.92.72.72.72.72.72.72.72.766661212121215.515.515.515.520202020Size (B parameters)
(a) Model Size
MonoCoderPolyCoderGPT-NeoGPT-JCodexStarCoderGPT-NeoX10101010202020203030303040404040505050503.513.513.513.512.332.332.332.333.693.693.693.692.822.822.822.822.552.552.552.551.711.711.711.712.372.372.372.373.163.163.163.166.296.296.296.299.969.969.969.9616.9216.9216.9216.9230.630.630.630.626.5126.5126.5126.5147.447.447.447.4Perplexity
(b) Perplexity (in green) and Normalized
-to-size Perplexity (in orange) for C
MonoCoderPolyCoderGPT-NeoGPT-JCodexStarCoderGPT-NeoX10101010202020203030303040404040505050503.693.693.693.692.992.992.992.992.872.872.872.872.472.472.472.471.951.951.951.952.012.012.012.012.322.322.322.323.323.323.323.328.078.078.078.077.757.757.757.7514.8214.8214.8214.8223.423.423.423.431.1631.1631.1631.1646.446.446.446.4Perplexity
(c) Perplexity (in green) and Normalized
-to-size Perplexity (in orange) for C++
Figure 2: Comparison of code language models based on their model size, perplexities, and normalized-to-size perplexities for C and C++. The results demonstrate that smaller models, such as MonoCoder, tend to have much better normalized perplexity scores (lower is better), indicating better performance relative to their size. Data for PolyCoder, GPT-Neo, GPT-J, Codex, StarCoder, and GPT-NeoX are taken from [24] and [33].
IV Local Semantics Elimination (LSE)